This is a piecewise-defined function I need help with.
If f (x)= { x^2 If x < 0
2. If x= 0
2x+1. If x> 0
Find (a) f(-2). (b) f (0). (c) f (2)
I hope you can help me understand the way to work these problems! Thank you!

Answer by flame8855(424)

a)
f(-2)
x = -2 < 0
f(x) = x^2 for x<0
f(-2) = 4
b)
if x = 0 , f(x) = 2
c)
f(2)
x=2 >0
f(x) = 2x+1 = 2(2)+1 = 5
f(2) = 5